# Homogenized C. elegans Neural Activity and Connectivity Data

### v2

@@ 2.3.1 Graph Tensor Data Processing
The data processing is handled by the ConnectomeBasePreprocessor class, which standardizes the neuron indices, processes edge attributes, and constructs the final graph tensors.
Table 4: Attributes in Graph Tensor Format. Attributes and their descriptions in the standardized graph tensor format. Includes neuron identities, edge connectivity types, synaptic weights, and spatial coordinates.
Symbol Description
$k$ Number of labeled neurons in a dataset
$l$ Number of labeled edges in a dataset
$N_{\text{pre}}^{k}$ Enumerated neuron identities for presynaptic neurons
$N_{\text{post}}^{k}$ Enumerated neuron identities for postsynaptic neurons
$T^{k}$ Connection type binary values (gap junction vs. chemical synapse)
$E$ Presynaptic neuron, postsynaptic neuron pairs which form edges
$G^{l}$ Gap junction weights for $l$ edges
$C^{l}$ Chemical synapse weights for $l$ edges
$x,y,z$ Cartesian coordinate $x,y,z$ position values for each neuron
1. Column Data Format.
Within each source connectome, presynaptic and postsynaptic neurons were specified along with chemical synapse and gap junction weight values. In the majority of the source connectome files, the format tabular for $k$ neurons with a column for the presynaptic neuron identity denoted by $N_{\text{pre}}^{k}$, a column for postsynaptic neuron identity denoted by $N_{\text{post}}^{k}$, connection type (gap junction vs. chemical synapse) denoted by $T^{k}$ with binary value elements (0 or 1), and the number of synapses, $N_{\text{weight}}^{k}$.
Table 5: Column Data Format for Connectome Dataset. Example of tabular data structure used in connectome datasets. Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters. The example shown is the first and last few rows of the OpenWorm [21] connectome source file
. Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
2. Matrix Connectome Data Format.
The exception to the tabular connectome source dataset format with distinct columns for presynaptic and postsynaptic neuron identities was the functional connectome from [14] and the connectome from [3]. These connectomes were in $N\times N$ matrix format where $N$ is the number of neurons. Neurons were arranged alphabetically. The $(i,j)$ element within the matrix represents the synaptic strength proportional to elicited response of surrounding stimulated neurons for the functional connectome from [14]. Similarly, the $(i,j)$ element within the matrix represents the number of synapses between neurons $i$ and $j$ for the connectome from [3]
3. Extraction.
Each connectome has a custom class created which inherits from the parent ConnectomeBasePreprocessor class. This was due to each connectome source being formatted differently from each other requiring custom loading and preprocessing from different file types in different formats.
Within each custom subclass, we implemented a preprocess function which was designed to open the source dataset respective file type and extract the relevant data to construct the final graph tensor object. Each preprocess function ends with the graph tensor object being saved in tensor file format via the save_graph_tensor() function call. Depending on the structure of the source dataset files, each custom preprocess function contained code responsible for opening and iterating through the different blocks of the file to extract relevant connectivity and edge weights data.
We used the torch_geometric library to instantiate a Data graph object which standardized the data structure used by the various connectome sources.
4. Graph Tensor Formatting
For each source connectome dataset, we have a list of neurons $N$ and $l$ edges $E\in\{N_{pre},N_{post}\}^{l}$. The list of edges between neurons which form the connectome are saved as the edge_index attribute of the graph tensor object. Additionally, synaptic weights for both gap junctions and chemical synapses were extracted from each source connectome dataset and saved as the edge_attr attribute of the graph tensor object. The gap junction weights, denoted by $G\in R^{l}$, and chemial synapse weights, denoted by $C\in R^{l}$, were appended together to produce the edge_attr tensor attributes. These weightages provide a standardized denotion of neuron influence to better construct graph conditioned, informed, and constrained neural activity models.
5. Cross-Dataset Pairing
The source connectome datasets did not contain all of the necessary information to produce the graph tensor object which was subsequently saved as a tensor file. Specifically, information such as the neuron type used by certain chemical synapse connections between neurons were not provided with every single source connectome dataset. As a result, we obtained connectivity information from each source dataset and paired that with neuron information associated with each chemical synapse edge. Additionally, other attributes were also shared across datasets which were not present in the original source datasets such as $x,y,z$ position of neurons. This cross-dataset pairing strategy allows for a more comprehensive and standardized graph tensor object format across different source datasets.
2.3.2 Consensus Connectome Analysis
Analysis across the different connectome source datasets was performed to extract commonalities between specific source connectome chemical synapse and gap junction connections. Specifically, we created a consensus connectome that, for every neuron-to-neuron chemical synapse or gap junction connection across all source datasets, selects the maximum weight value.
@@ 3 Discussion
In this paper we present a normalized and standardized dataset for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. One notable clarification is that while C. elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons. This shift in classification is supported by [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
@@
One of the primary limitations arises from the use of calcium fluorescence imaging as the method for measuring neural activity. Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6]. Calcium signals lag behind the actual electrical activity due to the slower dynamics of calcium ion concentration changes compared to the fast voltage changes associated with action potentials. This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost. Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that is critical for certain neural circuits [23, 20]. This highlights the need to complement calcium imaging with other techniques, such as voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity.
@@
Additionally, while the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity. This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching. The generation of the C. elegans connectome, as described by [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].
@@
On the computational side, while this dataset serves as a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using this dataset for machine learning models. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the neural dynamics data, affected by the limitations of calcium imaging, might bias models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional datasets (e.g., voltage imaging) to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. Graph neural networks (GNNs), for instance, are well-suited to handle the structural connectivity data, while recurrent neural networks (RNNs) and other time-series models can be applied to the neural activity data to capture temporal dependencies [9]. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the dataset’s open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
In conclusion, while this dataset offers significant opportunities for advancing our understanding of neural systems, its limitations should be carefully considered in both biological and computational contexts. Calcium imaging, structural connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used. As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C. elegans and the more complex neural architectures found in larger organisms [13, 12].
@@ 4 Conclusion
This unified dataset offers an accessible and standardized resource, streamlining access to data recorded under diverse experimental conditions and formats. The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data. Additionally, the full code repository containing the preprocessing code can be accessed on GitHub at worm-graph.
@@
The open-source nature of this resource aims to further contribute to both neuroscience and AI research, accelerating efforts to model not only C. elegans but also organisms with larger and more complex nervous systems. By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural dynamics modeling and the understanding of nervous systems across species.

### v3

@@ 2.3.2 Data Extraction and Preprocessing
The data extraction and preprocessing pipeline is implemented in the ConnectomeBasePreprocessor class, which standardizes neuron indices, processes edge attributes, and constructs the final graph tensor Data objects. Each source dataset is handled by a customized subclass since data extraction and preprocessing steps vary by dataset format:
  1. 1.
    Tabular Formats: Most datasets provide neuron pairs, connection types, and weights in a tabular structure. For example, the OpenWorm [22] dataset specifies presynaptic and postsynaptic neurons, connection types (chemical or gap junction), and weights. See Table 5 for an example.
  2. 2.
    Adjacency Matrix Formats: Some datasets use $D\times D$ adjacency matrices, with entries representing synapse counts ([3]) or functional connectivity values ([14]).
  3. 3.
    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not available from the source datasets. We obtain these from published research (e.g., neuron positions from [16], neurotransmitters and cell types from [27]) to supplement our graph structures.
Custom classes within the preprocessing pipeline handle dataset-specific formats, ensuring consistent graph tensor outputs.
2.3.3 Graph Tensor Format
To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with graph neural network (GNN) libraries. Each dataset is transformed into a torch_geometric [21] Data object with the following components:
  • edge_index: Adjacency list encoding $\mathcal{E}$.
  • edge_attr: Attribute vectors $\mathbf{a}_{ij}=[c,g,f,\dots]$ for edges in edge_index.
  • x: Node feature matrix $\mathbf{X}\in\mathbb{R}^{D\times L}$, initialized to zeros or populated with temporal slices of neural activity $\mathbf{X}^{k}[t:t+L]$ for some worm $k$. This data structure is the interface that allows integration of the neural activity data with connectivity data.
  • kwargs: Metadata and node attributes including neuron coordinates, class, type, and neurotransmitter.
This tensor format ensures compatibility with downstream ML models and facilitates integration of functional ($\mathbf{X}$) and structural ($\mathcal{E}$) data. Table 5 provides an overview of the standardized attributes.
2.3.4 Consensus Connectome
To address discrepancies in counts and weights of connections between datasets, we propose a consensus connectome that aggregates data across sources using majority voting. For each edge $e_{ij}$, the consensus weights are:
$c_{\text{consensus}}=\max\{c_{1},c_{2},\dots,c_{n}\},\quad g_{\text{consensus}}=\max\{g_{1},g_{2},\dots,g_{n}\},$
where $c_{i},g_{i}$ are the chemical and gap junction weights from $i$-th connectome source.
In order to quantify the uncertainty of this connectome, we also compute a weighted standard deviation $v$ of each edge across sources and include this as an additional attribute $\mathbf{a}=[c,g,...,v]$:
$v_{\text{consensus}}=\lambda\cdot\operatorname{std}\{c_{1},c_{2},\dots,c_{n}\}+(1-\lambda)\cdot\operatorname{std}\{g_{1},g_{2},\dots,g_{n}\}$
where $\lambda$ is a tuneable hyperparameter determining the relative importance of the chemical versus electrical synapse weights. We set $\lambda=0.9$ reflecting our knowledge that annotations of gap junctions are less certain than those of chemical synapses and that gap junctions are more sparse than chemical synapses in the connectome [3].
Having a single consensus connectome could used to model the average C. elegans nervous system, which has a stereotyped neurodevelopmental trajectory [35], or to quantify the inter-animal variability in network weights, which may reflect environmental influences. Figure 5 visualizes the consensus connectome of the 22 amphid chemosensory neurons [11].
Figure 3: Graph-Based Representation of the Connectome. Graph-based structure used for the C. elegans connectome data. Nodes represent neurons, with node attributes (e.g., position, class) and optional features. Directed edges capture connections, with edge attributes (e.g., chemical and gap junction weights).
@@ 3 Discussion
In this paper we present a normalized and standardized datasets for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code is composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
The homogenized neural and connectome datasets provide a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets. There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. We address most most of these in the next few paragraphs.
@@
One of the primary limitations of our neural activity dataset arises from the use of calcium fluorescence imaging as the method for measuring neural activity. Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6]. Calcium signals lag behind the actual electrical activity due to the slower dynamics of calcium ion concentration changes compared to the fast voltage changes associated with action potentials. This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost. Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that may be critical in certain neural circuits [33, 20]. This highlights the need to complement calcium-based measurements with other techniques, such as those utilizing voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity. However, calcium dynamics may be sufficient for the study of neural signaling in C. elegans as most of its neurons are nearly isopotential in the steady state [23].
@@
While the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and across the lifespan of a single individual [36]. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity measurements. This creates both an opportunity and a limitation. The opportunity arises from the fact that any robust relationship between structure and function to be discovered from this data must generalize across different age and phenotype-matched animals. The limitation is that this may preclude the development of individual-specific models which are desirable for understanding learning and memory, for example. While having the C. elegans connectome [35, 3] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons [14] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [1, 31].
@@
One important clarification about our work needs to made regarding our use of $300$ for the number of neurons in C. elegans, whereas readers may be more familiar the number $302$ from prior literature. We drop the bilateral pair of canal-associated neurons (CAN) from our datasets because they completely lack chemical synapses and are now widely classified as ‘end-organs’ rather than neurons [17, 3]. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the used of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. For instance, graph neural networks (GNNs) [34] are well-suited to handle the structural connectivity data, while time-series models like structured state space models [25] can be applied to the neural activity data to capture temporal dependencies. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the datasets’ open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
In conclusion, while our homogenized datasets offer significant opportunities for advancing our understanding of small neural systems, its limitations should be carefully considered in both biological and computational contexts. Calcium imaging, noisy connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used. As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C. elegans and the more complex neural architectures found in larger organisms.
@@ 4 Conclusion
This unified dataset offers an accessible and standardized resource, streamlining access to data recorded under diverse experimental conditions and formats. The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data. Additionally, the full code repository containing the preprocessing code can be accessed on GitHub.
@@
The open-source nature of this resource aims to further contribute to both neuroscience and AI research, accelerating efforts to model not only C. elegans but also organisms with larger and more complex nervous systems. By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural structure-function modeling and the discovery of principles of neural computation across species.